Among the Natural no.s {1,2,3,.....23,24,25},in how many ways we can draw 11 no.s so that no two of them are consecutive?



Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Borah, 

first let there be a gap of one number between every two consecutive selections. when this is done, we have fixed 11*2 -1 =21 consecutive numbers. each of the remaining 4 numbers can be inserted into any one of 12 spaces. thus the number of ways=12*12*12*12=124

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras